# Scalable and consistent embedding of probability measures into Hilbert spaces via measure quantization

### v1

@@ 4.4 Earth image dataset
Table 2: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on ${K}$-LOT and $K$-KME with both quantization methods.
$\overline{K}$-LOT/$\tilde{K}$-LOT $\overline{K}$-KME/$\tilde{K}$-KME
$K$ Accuracy Time (s) Accuracy Time
16 88/88 15/228 76/68 219/732
32 89/89 17/249 67/67 2792/2247
64 89/88 20/305 68/68 7958/9246
128 88/88 32/390 67/66 31765/32287
We perform similar experiments on a set of images provided by the Airbus company. The dataset consists in $N=1000$ images of size $128\times 128$ captured by a SPOT satellite. The images are divided into two categories : those with the presence of a wind turbine and those without, see Figures 7(a) and 7(b) in Appendix C. Each image is viewed as a discrete probability distribution on the RGB space, that is each pixel is represented by a point in $\mathbb{R}^{3}$. The size $N$ of the dataset as well as the number of pixels ($m_{i}=128^{2}$) prevents from directly computing either LOT or KME. We therefore carry out supervised classification from both embeddings and both quantization methods. For $K\in\{16,32,64,128\}$, we implement $K$-LOT with reference measure the uniform measure on $[0,1]^{3}$ sampled on $m_{0}=1000$ points, and $K$-KME with the Gaussian kernel with bandwidth parameter $\sigma=100$. After the embeddings, we perform PCA and retain only the first 10 components, on which we train an LDA classifier using 75% of the data and test it on the remaining 25%. We obtain the accuracies and execution times displayed in Table 2. Both quantization methods achieve similar accuracy results while mean-measure quantization is approximately ten times faster than the quantization of each measure.
@@ Appendix A Proofs of the main results
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$ the set of distinct points and consider the Voronoï partition
V~x1=Vx1 and
V~xk
@@
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=\sum_{k=1}^{K}\int_{\widetilde{V}_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu(y)$
@@
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$. From the dual formulation of the Kantorovich problem (see e.g. [44]), we have that, for any $a\in\Sigma_{K}$,
$W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\sup\limits_{\beta\in\mathbb{R}^{K}}\sum\limits_{k=1}^{K}a_{k}\beta_{k}+\int_{\mathbb{R}^{d}}\left(\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}-\beta_{k}\}\right)\mathrm{d}\mu(y)$ (24)
@@
where the above inequality is obtain by taking $\beta_{k}=0$ for all $1\leq k\leq K$. Then, since $X=(x_{1},\cdots,x_{K})\in F_{K}$, one has that $(\widetilde{V}_{x_{k}})_{1\leq k\leq K}$ is a partition of $\mathbb{R}^{d}$, and we may define
@@
that is a mapping from $\mathbb{R}^{d}$ to $X$. Introducing the probability measure $\mu_{K}=\sum\limits_{k=1}^{K}\mu(\widetilde{V}_{x_{k}})\delta_{x_{k}},$ it is not difficult to see that $T_{K\#}\mu=\mu_{K}$ where the notation $T_{\#}\mu$ denotes the push-forward of a measure $\mu$ by the mapping $T$. Now, we let $\pi_{K}=(\mathrm{id}\times T_{K})_{\#}\mu$ that obviously belongs to the set of transport plans $\Pi(\mu,\mu_{K})$. From the definition of the Voronoï partition $(\widetilde{V}_{x_{k}})_{1\leq k\leq K}$, one can than check that, for any $y\in\mathbb{R}^{d}$, (see e.g. [35])
@@
Inserting the above equality into (24), we thus have that, for any $a\in\Sigma_{K}$,
@@
$\min\limits_{a^{(i)}\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}=\sum_{k=1}^{K}\int_{\widetilde{V}_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y),$
@@
$\min\limits_{a\in(\Sigma_{K})^{N}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\int_{\widetilde{V}_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y)$
$=$ $\sum\limits_{k=1}^{K}\int_{\widetilde{V}_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\overline{\mu}(y)$
$=$ $W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}$
@@
$\min\limits_{X\in F_{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}=\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(\widetilde{V}_{x_{k}})\delta_{x_{k}}\Bigr{)}$
@@
Proof of Proposition 3.4.
For a fixed $N\geq 1$, since $\mathbb{P}^{N}$, ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ are discrete uniform measures of the same size, one can actually restrict the search of an optimal plan in (4) as that of finding an optimal permutation $\sigma\in\mathrm{Perm}(N)$ see e.g. [44] in the following sense:
@@
$\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\widetilde{\mathbb{P}}^{N}_{K})$ $=\min\limits_{\sigma\in\mathrm{Perm}(N)}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(\sigma(i))}_{K})$ (25)
However, for $\tilde{\nu}^{(i)}_{K}$, defined in (11), it follows by the definition of optimal quantization of $\mu^{(i)}$ that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(j)}_{K}).$ (26)
Now for $\bar{\nu}^{(i)}_{K}$, defined in (14). Since $\bar{\nu}^{(i)}_{K}$ corresponds to the discrete probability measure supported on $\bar{X}$ that best approximates $\mu^{(i)}$. In other words, $W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\sum_{k=1}^{K}a_{k}\delta_{\bar{x}_{k}})$ for any weight vector $a\in\Sigma_{K}$. In particular, we have that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(j)}_{K}).$ (27)
Using Inequalities (26) and (27), it is then easy to see that in both cases, the optimal permutation minimizing (25) is the identity $\sigma(i)=i$ for all $1\leq i\leq N$, and that we have
@@
$\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})=W_{2}^{2}\left(\sum\limits_{k=1}^{K}\overline{\mu}(\tilde{V}_{\bar{x}_{k}})\delta_{\bar{x}_{k}},\overline{\mu}\right)=\varepsilon_{K}$
@@
As $\mathrm{MMD}$ is an integral probability metric [20][Lemma 4], it is bounded by the $1$-Wasserstein distance $W_{1}$ defined by :
$W_{1}(\rho,\mu)=\min\limits_{\pi\in\Pi(\rho,\mu)}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|\mathrm{d}\pi(x,y).$ (28)
Since $W_{1}\leq W_{2}$ [40][Chapter 5], we thus have that $\mathrm{MMD}\leq W_{2}$. Therefore,
$\frac{1}{N}\sum_{i=1}^{N}\text{MMD}^{2}(\mu^{(i)},\nu^{(i)}_{K})\leq\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$ (29)
@@
For a fixed $N$, and thanks to Proposition 3.4, we have that the sequence of probability measures $(\mathbb{P}^{N}_{K})_{K\geq 1}$ such that $\mathbb{P}^{N}_{K}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\subset\mathcal{P}(\mathcal{P}(\mathcal{X}))$ converges towards $\mathbb{P}^{N}$, that is $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P}_{N})\overset{K\rightarrow\infty}{\longrightarrow}0$. Additionally, the Wasserstein barycenter of $\mathbb{P}^{N}$ is unique (Proposition 3.5 in [1]) since at least one of the probability measures $\mu^{(i)},1\leq i\leq N$ is a.c. by hypothesis. Therefore, using [27][Theorem 3], we immediately obtain that the sequence of barycenters of $(\mathbb{P}^{N}_{K})_{K\geq 1}$ converges towards the barycenter of $\mathbb{P}^{N}$ in $W_{2}$ distance.
@@
$(a+b+c)^{2}\leq(2+\lambda)a^{2}+(2+\lambda)c^{2}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}c^{2}$ (30)
Squaring the triangle inequality and using (30), this yields to:
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq(2+\lambda)W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}W_{2}^{2}(\mu^{(i)},\mu^{(j)})+(2+\lambda)W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})\bigr{)}$ (31)
We now sum inequality (31) over all $1\leq i,j\leq N$, and divide by $N^{2}$:
$\mathrm{SS}(\nu_{K})$ $=\frac{1}{N^{2}}\sum\limits_{i=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})\leq\frac{2}{N}(2+\lambda)\sum\limits_{i=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\frac{1}{N^{2}}\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\sum\limits_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)})$
Hence, by Proposition 3.4, we obtain that $\mathrm{SS}(\nu_{K})\leq(4+2\lambda)\varepsilon_{K}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\mathrm{SS}(\mu)$, which concludes the proof. ∎

### v3

@@ 5.4 Earth image dataset
Table 2: Earth image dataset. LDA classification accuracy on the Airbus dataset after 10-component PCA on the methods with $K=64$.
Method Accuracy Time (s)
$\overline{K}$-LOT 89 20
$\tilde{K}$-LOT 88 305
Random subset of size $K$ + LOT 77 35
Random subset of size $K$ + KME 70 8011
$\overline{K}$-KME 68 7958
$\tilde{K}$-KME 68 9246
$K$-Nys-KME 65 189
We perform similar experiments on a set of images provided by the Airbus company. The dataset consists in $N=1000$ images of size $128\times 128$ captured by a SPOT satellite. The images are divided into two categories : those with the presence of a wind turbine and those without, see Figures 8(a) and 8(b) in Appendix C. Each image is viewed as a discrete probability distribution on the RGB space, that is each pixel is represented by a point in $\mathbb{R}^{3}$. The size $N$ of the dataset as well as the number of pixels ($m_{i}=128^{2}$) prevents from directly computing either LOT or KME. We therefore carry out supervised classification from both embeddings and both quantization methods. We implement $K$-LOT with reference measure the uniform measure on $[0,1]^{3}$ sampled on $m_{0}=1000$ points, and $K$-KME with the Gaussian kernel with bandwidth parameter $\sigma=100$. After the embeddings, we perform PCA and retain only the first 10 components, on which we train an LDA classifier using 75% of the data and test it on the remaining 25%. We obtain the accuracies and execution times displayed in Table 2 for $K=64$, with additional experiments in Appendix C.3 for different values of $K$. Both quantization methods achieve similar accuracy results while mean-measure quantization is approximately ten times faster than the quantization of each measure. We continue to observe that LOT outperforms KME, and that applying quantization to KME yields results comparable to those of other KME-based approaches.
@@ 6 Conclusion
One limitation of our quantization approach is its sensitivity to the curse of dimensionality‚ a challenge common to many statistical problems in optimal transport. In future works, one could bypass the dependance on the dimension by relying on some notion of intrinsic dimension, as studied in [53].
7 Acknowledgments
This work benefited from financial support from the French government managed by the National Agency for Research (ANR) under the France 2030 program, with the reference ANR-23-PEIA-0004. The authors also gratefully acknowledge another financial support from the ANR (MaSDOL grant ANR-19-CE23-0017).
@@ Appendix A Proofs of the main results
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$ in (2.7) a vector of distinct points and consider the Voronoï partition
Ux1=Vx1 and
Uxk
@@
$\min\limits_{a\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu(y)$
@@
Let $X=(x_{1},\cdots,x_{K})\in F_{K}$. From the dual formulation of the Kantorovich problem [40][Equation 5.7], we have that, for any $a\in\Sigma_{K}$,
$W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}a_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\sup\limits_{\beta\in\mathbb{R}^{K}}\sum\limits_{k=1}^{K}\beta_{k}a_{k}+\int_{\mathbb{R}^{d}}\beta^{c}(y)\mathrm{d}\mu(y)$ (A.1)
$=$ $\sup\limits_{\beta\in\mathbb{R}^{K}}\sum\limits_{k=1}^{K}a_{k}\beta_{k}+\int_{\mathbb{R}^{d}}\left(\min_{1\leq k\leq K}\{\|x_{k}-y\|^{2}-\beta_{k}\}\right)\mathrm{d}\mu(y)$
@@
where the above inequality is obtain by taking $\beta_{k}=0$ for all $1\leq k\leq K$. Then, since $X=(x_{1},\cdots,x_{K})\in F_{K}$, one has that $(U_{x_{k}})_{1\leq k\leq K}$ is a partition of $\mathbb{R}^{d}$, and we may define
@@
that is a mapping from $\mathbb{R}^{d}$ to $X$. Introducing the probability measure $\mu_{K}=\sum\limits_{k=1}^{K}\mu(U_{x_{k}})\delta_{x_{k}},$ it is not difficult to see that $T_{K\#}\mu=\mu_{K}$ where the notation $T_{\#}\mu$ denotes the push-forward of a measure $\mu$ by the mapping $T$. Now, we let $\pi_{K}=(\mathrm{id}\times T_{K})_{\#}\mu$ that obviously belongs to the set of transport plans $\Pi(\mu,\mu_{K})$. From the definition of the Voronoï partition $(U_{x_{k}})_{1\leq k\leq K}$, one can than check that, for any $y\in\mathbb{R}^{d}$, (see e.g. [39])
@@
Inserting the above equality into (A.1), we thus have that, for any $a\in\Sigma_{K}$,
@@
$\min\limits_{a^{(i)}\in\Sigma_{K}}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}=W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y),$
@@
$\min\limits_{a\in(\Sigma_{K})^{N}}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu^{(i)},\sum\limits_{k=1}^{K}a^{(i)}_{k}\delta_{x_{k}}\Bigr{)}$ $=$ $\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}\Bigl{(}\mu,\sum\limits_{k=1}^{K}\mu^{(i)}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\mu^{(i)}(y)$
$=$ $\sum\limits_{k=1}^{K}\int_{U_{x_{k}}}\|x_{k}-y\|^{2}\mathrm{d}\overline{\mu}(y)$
$=$ $W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
@@
$\min\limits_{X\in F_{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}=\min\limits_{X\in(\mathbb{R}^{d})^{K}}W_{2}^{2}\Bigl{(}\overline{\mu},\sum\limits_{k=1}^{K}\overline{\mu}(U_{x_{k}})\delta_{x_{k}}\Bigr{)}$
@@
Proof of Theorem 3.4.
For a fixed $N\geq 1$, since $\mathbb{P}^{N}$, ¯$\overline{\mathbb{P}}^{N}_{K}$ and ~$\widetilde{\mathbb{P}}^{N}_{K}$ are discrete uniform measures of the same size, one can actually restrict the optimization set in (2.3) to permutations $\sigma\in\mathrm{Perm}(N)$ [40][Equation 2.2] in the following sense:
@@
$\mathcal{W}_{2}^{2}(\mathbb{P}^{N},\widetilde{\mathbb{P}}^{N}_{K})$ $=\min\limits_{\sigma\in\mathrm{Perm}(N)}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(\sigma(i))}_{K})$ (A.2)
However, for $\tilde{\nu}^{(i)}_{K}$, defined in (3.1), it follows by the definition of optimal quantization of $\mu^{(i)}$ that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\tilde{\nu}^{(j)}_{K}).$ (A.3)
Similarly, $\bar{\nu}^{(i)}_{K}$ defined in (3.4) corresponds to the discrete probability measure supported on $\bar{X}$ that best approximates $\mu^{(i)}$. In other words, $W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\sum_{k=1}^{K}a_{k}\delta_{\bar{x}_{k}})$ for any weight vector $a\in\Sigma_{K}$. In particular, we have that, for any $1\leq j\leq N$,
$W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})\leq W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(j)}_{K}).$ (A.4)
Using Inequalities (A.3) and (A.4), it is then easy to see that in both cases, the optimal permutation minimizing (A.2) is the identity $\sigma(i)=i$ for all $1\leq i\leq N$, and that we have
@@
$\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\bar{\nu}^{(i)}_{K})=W_{2}^{2}\left(\sum\limits_{k=1}^{K}\overline{\mu}(U_{\bar{x}_{k}})\delta_{\bar{x}_{k}},\overline{\mu}\right)=\varepsilon_{K}(\overline{\mu}),$
@@
We have that $(\mathcal{X},\|\cdot\|)$ is a complete separable metric space, $k$ is positive definite and verifies $\forall x,y\in\mathcal{X},k(x,x)+k(y,y)-2k(x,y)\leq C^{2}\|x-y\|^{2}$ for a constant $C>0$, and $\mu^{(i)}$ and $\nu^{(i)}$ have finite 2-moments as they are supported on a compact set $\mathcal{X}\subset\mathbb{R}^{d}$. We can then use Proposition 2 of [50], which proves that under these assumptions,
$\mathrm{MMD}(\mu^{(i)},\nu_{K}^{(i)})\leq CW_{2}(\mu^{(i)},\nu_{K}^{(i)}).$
We can directly conclude with Theorem 3.4 that:
$\frac{1}{N}\sum_{i=1}^{N}\text{MMD}^{2}(\mu^{(i)},\nu^{(i)}_{K})$$\leq$ $\frac{1}{N}\sum_{i=1}^{N}C^{2}W_{2}^{2}(\mu^{(i)},\nu^{(i)}_{K})$
$=$ $C^{2}\varepsilon_{K}\overset{K\rightarrow\infty}{\longrightarrow}0.$
Proof of Theorem 3.9.
The proof of Theorem 3.9 relies on [10][Theorem 5.2], that we recall below for completeness.
Theorem A.2 (Theorem 5.2 from [10]).
Let $\mu\in\mathcal{P}(\mathcal{X})$ with absolutely continuous part $f\neq 0$. For $\tau>0$, all sufficiently large values of $m$ and $K=C\cdot M(f)\cdot m^{d/(2d+4)}$, it holds that
$W_{2}(\mu,\hat{\nu}_{K})\leq C\cdot m(f)\cdot m^{-1/(2d+4)\cdot\tau}~{}\text{with probability}~{}1-e^{-\tau^{2}},$
where $\hat{\nu}_{K}$ is the $K$-points optimal quantization of the empirical measure $\hat{\mu}=\frac{1}{m}\sum_{j=1}^{m}\delta_{x_{j}}$ associated to $\mu$ supported on $m$ points and $C$ is a constant that only depends on $d$.
Let $\hat{\nu}^{(i)}_{K^{(i)}}$ be the $K^{(i)}$-points optimal quantization of $\hat{\mu}^{(i)}$, the empirical version of $\mu^{(i)}$ supported on $m_{i}$ points. Adapting Theorem A.2 in this context gives that for $\tau>0$, for sufficiently large $m_{i}$ and letting $K^{(i)}=C\cdot m(f^{(i)})\cdot m_{i}^{d/(2d+4)}$, it holds that
$W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}})\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}~{}\text{with~{} probability}~{}1-e^{-\tau^{2}},$ (A.5)
In other words, Equation (A.5) reads
$\mathbf{Pr}\Bigl{[}W_{2}^{2}\bigl{(}\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-e^{-\tau^{2}}$
If we denote $A_{i}$ the event $\{W_{2}^{2}\bigl{(}\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}>C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\}$, we obtain that
$\mathbf{Pr}[A_{i}]\leq e^{-\tau^{2}}\Longrightarrow\mathbf{Pr}\Bigl{[}\cup_{i=1}^{N}A_{i}\Bigr{]}\leq\sum\limits_{i=1}^{N}e^{-\tau^{2}}=Ne^{-\tau^{2}},$
which can be rewritten as
$\mathbf{Pr}$ $\Bigl{[}\exists i,W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}>C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\leq Ne^{-\tau^{2}}$
$\Longrightarrow$ $\mathbf{Pr}$ $\Bigl{[}\forall i,W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
$\Longrightarrow$ $\mathbf{Pr}$ $\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K^{(i)}}\bigr{)}\leq\frac{1}{N}\sum\limits_{i=1}^{N}C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
Now, taking $K=\max\limits_{1\leq i\leq N}K^{(i)}$, one has $\forall 1\leq i\leq N$, $W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K})\leq W^{2}_{2}(\mu^{(i)},\hat{\nu}^{(i)}_{K^{(i)}})$. Hence, we finally obtain that
$\mathbf{Pr}\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K}\bigr{)}\leq\frac{1}{N}\sum\limits_{i=1}^{N}C^{2}\cdot m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\cdot\tau^{2}\Bigr{]}\geq 1-Ne^{-\tau^{2}}$
Now taking $\tau=\sqrt{\delta^{2}+\log(N)}$, one has that
$\mathbf{Pr}\Bigl{[}\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}\bigl{(}\mu^{(i)},~{}\hat{\nu}^{(i)}_{K}\bigr{)}\leq C^{2}\frac{\delta^{2}+\log(N)}{N}\sum\limits_{i=1}^{N}m(f^{(i)})^{2}\cdot m_{i}^{-1/(d+2)}\Bigr{]}\geq 1-e^{-\delta^{2}}.$
We can conclude the proof of Theorem 3.9 by observing that $\mathcal{W}_{2}^{2}(\hat{\mathbb{P}}_{K}^{N},\mathbb{P}^{N})\leq\frac{1}{N}\sum\limits_{i=1}^{N}W_{2}^{2}(\mu^{(i)},\hat{\nu}_{K}^{(i)})$.
@@
For a fixed $N$, and thanks to Theorem 3.4, we have that the sequence of probability measures $(\mathbb{P}^{N}_{K})_{K\geq 1}$ such that $\mathbb{P}^{N}_{K}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\nu^{(i)}_{K}}\subset\mathcal{P}(\mathcal{P}(\mathcal{X}))$ converges towards $\mathbb{P}^{N}$, that is $\mathcal{W}_{2}(\mathbb{P}^{N}_{K},\mathbb{P}_{N})\overset{K\rightarrow\infty}{\longrightarrow}0$. Additionally, the Wasserstein barycenter of $\mathbb{P}^{N}$ is unique (Proposition 3.5 in [1]) since at least one of the probability measures $\mu^{(i)},1\leq i\leq N$ is a.c. by hypothesis. Therefore, using [31][Theorem 3], we immediately obtain that the sequence of barycenters of $(\mathbb{P}^{N}_{K})_{K\geq 1}$ converges towards the barycenter of $\mathbb{P}^{N}$ in the $W_{2}$ distance.
@@
$(a+b+c)^{2}\leq(2+\lambda)a^{2}+(2+\lambda)c^{2}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}b^{2}$ (A.6)
Squaring the triangle inequality and using (A.6) yields to:
$W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq(2+\lambda)W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}W_{2}^{2}(\mu^{(i)},\mu^{(j)})+(2+\lambda)W_{2}^{2}(\mu^{(j)},\nu^{(j)}_{K})$ (A.7)
We now sum inequality (A.7) over all $1\leq i,j\leq N$, and divide by $N^{2}$:
$\mathrm{SS}(\nu_{K})$ $=\frac{1}{N^{2}}\sum\limits_{i,j=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\nu^{(j)}_{K})\leq\frac{2}{N}(2+\lambda)\sum\limits_{i=1}^{N}W_{2}^{2}(\nu^{(i)}_{K},\mu^{(i)})+\frac{1}{N^{2}}\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\sum\limits_{i,j=1}^{N}W_{2}^{2}(\mu^{(i)},\mu^{(j)})$
Hence, by Theorem 3.4, we obtain that $\mathrm{SS}(\nu_{K})\leq(4+2\lambda)\varepsilon_{K}+\Bigl{(}1+\frac{2}{\lambda}\Bigr{)}\mathrm{SS}(\mu)$, which concludes the proof. ∎